OpenStudy (anonymous):

The width of a rectangle is 5 cm less than half its length. Its area is 72cm squared. Find the length and the width of the rectangle

myininaya (myininaya):

W=1/2L-5 A=LW=72

myininaya (myininaya):

use these two equations

OpenStudy (anonymous):

but how do i do it?

myininaya (myininaya):

plug W=1/2L-5 into LW=72 and solve for L first

OpenStudy (anonymous):

ok i got the answer can you tell me so i can check to make sure it is correct?

myininaya (myininaya):

just plug 1/2L-5 where you see W in LW=72 like this L(1/2L-5)=72 now disbribute 1/2L^2-5L=72 subtract 72 on both sides 1/2L^2-5L-72=0 now go ahead in get rid of that nasty coefficient on L^2 by multiplying both sides by 2 L^2-10L-144=0 (L-18)(L+8)=0 L=18 or L=-8 negative length is crazy talk so we have L=18 now you can plug this value into W=1/2L-5 to find W